### v1

# Tracing Multilingual Representations in LLMs with Cross-Layer Transcoders

Abstract
Multilingual Large Language Models (LLMs) can process many languages, yet how they internally represent this diversity remains unclear. Do they form shared multilingual representations with language-specific decoding, and if so, why does performance still favor the dominant training language? To address this, we train a series of LLMs on different mixtures of multilingual data and analyze their internal mechanisms using cross-layer transcoders (CLT) and attribution graphs. Our results provide strong evidence for pivot language representations: the model employs nearly identical representations across languages, while language-specific decoding emerges in later layers. Attribution analyses reveal that decoding relies in part on a small set of high-frequency language features in the final layers, which linearly read out language identity from the first layers in the model. By intervening on these features, we can suppress one language and substitute another in the model’s outputs. Finally, we study how the dominant training language influences these mechanisms across attribution graphs and decoding pathways. We argue that understanding this pivot-language mechanism is crucial for improving multilingual alignment in LLMs.
Tracing Multilingual Representations in LLMs with
Cross-Layer Transcoders
Abir Harrasse1,2thanks: Equal contribution. First author order determined by coin flip. The first author’s work was done during her internship at the Max Planck Institute for Intelligent Systems., Florent Draye1111Models: https://huggingface.co/collections/CausalNLP/multilingual-tinystories-6862b6562414eb84d183f82a and https://huggingface.co/collections/CausalNLP/multilingual-gpt2-models-684ad70e5fb3c84962306af3, Zhijing Jin1,3,4, Bernhard Schölkopf1,5 1Max Planck Institute for Intelligent Systems, Tübingen, Germany 2Mohammed VI Polytechnic University, Morocco 3University of Toronto, Canada  4Vector Institute, Canada 5ELLIS Institute, Tübingen, Germany abir.harrasse@emines.um6p.ma fdraye@tue.mpg.de zjin@cs.toronto.edu
Figure 1: Overview of our study. We analyze the multilingual behavior of LLMs using attribution graphs derived from cross-layer transcoders (CLTs)Ameisen et al. (2025). These graphs provide fine-grained feature representations and capture their interactions more effectively than methods used in prior studies, such as Logit Lens.
1 Introduction
The inclusion of diverse languages in the large-scale training data of large language models (LLMs) Grattafiori et al. (2024); Team et al. (2024); Brown et al. (2020); Chowdhery et al. (2022) has led to remarkable multilingual capabilities Shi et al. (2022); Conneau et al. (2020a); Workshop et al. (2023) However, the underlying mechanisms driving their multilingual behavior remain poorly understood. For example, do LLMs form shared multilingual representations in their internal layers, so, why does performance still favor the dominant training language, such as English? Prior mechanistic investigations have yielded conflicting findings about multilingual processing. While some work maps inputs into English-aligned latent spaces and suggests that LLMs process non-English inputs through English-biased representations Wendler et al. (2024); Schut et al. (2025), other studies find evidence of shared cross-lingual grammatical structures encoded in common feature directions Brinkmann et al. (2025). Moreover, all existing approaches rely primarily on static probing methods that cannot capture the dynamic transformation of linguistic information across model layers.
To address the aforementioned limitations, we derive attribution graphs from our self-trained cross-layer transcoders Ameisen et al. (2025), a recent advance in mechanistic interpretability, which can track feature interactions across layers and provide the first mechanistic account of how LLMs transition between shared and language-specific processing modes. Our results suggest that all languages have a shared representation in LLMs’ internal layers, which we call as a “pivot language.” We also discover language-specific decoding in layer layers to handle their multilingual output generation.
Depending on our CLT-driven insights, we conduct two causal studies. First, we are able to change the output language by intervening on a small set of high-frequency language features in the final layers decoding from the pivot language to the target language. Further, we investigate the causal effect of training data composition by varying English-to-multilingual ratios. We analyze our re-trained models on different language ratios, and reveal how language distribution during pretraining shapes internal multilingual representations.
Our contributions are as follows:
  1. 1.
    We present comprehensive evidence for a pivot language in multilingual LLMs: models form shared multilingual spaces and use consistent mechanisms to process equivalent queries across languages. This holds across training mixtures, suggesting that pivot languages emerge independently of English dominance, which challenges the view that “models think in English”.
  2. 2.
    We show that language decoding primarily depends on a small set of high-frequency language features in the late layers that read of information from early layer features and tokens embeddings.
  3. 3.
    We identify and characterize language-specific failure modes, where models succeed in English but fail on equivalent inputs in other languages, linking these disparities to underlying mechanisms.
  4. 4.
    We release eight multilingual models [^1] Models: https://huggingface.co/collections/CausalNLP/multilingual-tinystories-6862b6562414eb84d183f82a and https://huggingface.co/collections/CausalNLP/multilingual-gpt2-models-684ad70e5fb3c84962306af3 (two sizes × four data mixtures) and their corresponding cross-layer transcoders (CLTs) [^2] CLTs: https://huggingface.co/flodraye, trained on balanced language distributions to support cross-lingual analysis and transfer.
2 Related Work
2.1 Multilingual Language Models
Multilingual language models have evolved from early architectures like mBERT Devlin et al. (2019) to sophisticated models such as XLM-R Conneau et al. (2020b), LLaMA Touvron et al. (2023), and Gemma Team et al. (2024). Despite their impressive capabilities, systematic performance disparities persist across languages Pires et al. (2019); Hu et al. (2020). English consistently outperforms other languages, particularly for low-resource languages and complex reasoning tasks Lauscher et al. (2020); Shi et al. (2022). Benchmarks like XTREME Hu et al. (2020) and XNLI Conneau et al. (2020b) have documented these gaps across language families. The English-dominant training data composition has been hypothesized as a primary factor Bender et al. (2021), though the underlying mechanisms driving these disparities remain poorly understood.
2.2 Internal Representations and the Pivot Hypothesis
Recent investigations have revealed that multilingual LLMs process non-English inputs through English-aligned latent representations. Wendler et al. Wendler et al. (2024) used logit lens analysis to show that intermediate layers predict English tokens before target-language tokens. Similarly, Schut et al. Schut et al. (2025) demonstrated systematic bias toward English-like representations regardless of input language. These findings suggest models internally translate to English, perform reasoning, then translate back, a mechanism that has been leveraged through embedding alignment Lample and Conneau (2019); Artetxe et al. (2018), translation-based fine-tuning Zhu et al. (2024), and romanization strategies for non-Latin scripts Saji et al. (2025).
Prompting strategies also exploit this English-pivot behavior, including direct prompt translation Shi et al. (2022); Ahuja et al. (2023); Etxaniz et al. (2023) and English chain-of-thought reasoning Wei et al. (2022); Vatsal et al. (2025). While some grammatical features exhibit universal representations across languages Brinkmann et al. (2025), lexical and semantic processing appears more language-specific Chi et al. (2020).
These observations align with the recently proposed “Platonic hypothesis” Huh et al. (2024), which suggests that models converge toward abstract, unified representations capturing conceptual essences beyond surface forms. The pivot-language behavior, where models develop language-agnostic representations before language-specific decoding, provides empirical support for this hypothesis in multilingual settings.
2.3 Mechanistic Interpretability in LLMs
Mechanistic interpretability seeks to understand transformer internals through circuit analysis Olah et al. (2020); Elhage et al. (2021). Early work identified specific circuits for behaviors like indirect object identification Wang et al. (2022) and induction heads Olsson et al. (2022). Sparse Autoencoders (SAEs) decompose representations into interpretable features Cunningham et al. (2023); Templeton et al. (2024), while transcoders directly model MLP outputs rather than autoencoding activations Dunefsky et al. (2024).
Cross-Layer Transcoders (CLTs) advance this approach by assigning distinct decoder matrices per downstream layer, significantly simplifying feature graphs and enabling tractable circuit analysis Ameisen et al. (2025); Hanna et al. (2025). However, CLT applications remain limited—primarily Anthropic’s foundational work Lindsey et al. (2025b), a “greater-than” mechanism study Merullo et al. (2025), and recent open-source extensions Lindsey et al. (2025a). This scarcity reflects the considerable challenges in training CLTs at scale.
For multilingual models, mechanistic analyses are particularly sparse. Beyond Wendler et al.’s logit lens study Wendler et al. (2024) and Anthropic’s case study Lindsey et al. (2025b), our work represents the first comprehensive CLT-based investigation of multilingual processing. Crucially, ensuring fair cross-language comparisons requires training CLTs on balanced multilingual distributions to avoid feature bias toward dominant languages.
3 Experimental Setup and Methods
3.1 Mixture of multilingual LLMs
We study the multilingual behavior of LLMs across five languages: English, German, French, Arabic, and Chinese. To investigate how training data composition affects internal representations, we train four GPT-2 style models (12 layers, $\sim$177.6M parameters) on a realistic web mixture (OpenWebText + FineWeb2), and four tiny-stories style models (4 layers and $\sim$68.5M parameters) on a controlled narrative dataset (TinyStories translated into all five languages). Finally, we include an off-the-shelf LLaMA 1B model for comparison.
To facilitate our training data manipulation study, we investigate training data settings where all languages are evenly distributed vs those with one dominant language. To investigate the effect of a dominant language, we vary the English proportion in the data from dominant (90%) to balanced (20%), with the remaining data evenly split across the other four languages. All models use debiased tokenizers trained on evenly distributed five-language data, so that differences in internal representations reflect the training distribution rather than tokenization artifacts. Figure 1 illustrates the training configurations.
3.2 Cross-Layer Transcoders (CLTs)
To probe internal multilingual mechanisms, we employ Cross-Layer Transcoders (CLTs), which map activations between layers to reveal how semantic representations evolve during processing. A cross-layer transcoder consists of neurons (features) divided into encoder and decoder components.
Formally, to run a cross-layer transcoder, let $\mathbf{h}_{\ell}\in\mathbb{R}^{d_{\text{model}}}$ be the input to the MLP at layer $\ell$ for single token position. We define
$\mathbf{z}_{\ell}=\text{ReLU}(\mathbf{W}_{\text{enc}}^{\ell}\mathbf{h}_{\ell}+\mathbf{b}_{\text{enc}}^{\ell})\in\mathbb{R}^{d_{\text{features}}},$ (1)
where $\mathbf{W}_{\text{enc}}^{\ell}\in\mathbb{R}^{d_{\text{features}}\times d_{\text{model}}}$ is the encoder weight matrix and $\mathbf{b}_{\text{enc}}^{\ell}\in\mathbb{R}^{d_{\text{features}}}$ is the encoder bias. The CLTs then reconstructs the MLP output at layer $\ell^{\prime}$ as
$\hat{\mathbf{m}}_{\ell^{\prime}}=\sum_{\ell\leq\ell^{\prime}}\mathbf{W}_{\text{dec}}^{\ell\to\ell^{\prime}}\mathbf{z}_{\ell}+\mathbf{b}_{\text{dec}}^{\ell^{\prime}},$ (2)
where $\mathbf{W}_{\text{dec}}^{\ell\to\ell^{\prime}}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{features}}}$ is the decoder matrix from layer $\ell$ to $\ell^{\prime}$.
Following Anthropic guidelines Ameisen et al. (2025), the final training objective is
$\mathcal{L}=$ $\underbrace{\sum_{\ell^{\prime}}\left\|\hat{\mathbf{m}}_{\ell^{\prime}}-\mathbf{m}_{\ell^{\prime}}\right\|_{2}^{2}}_{\text{MSE reconstruction}}$
$+\lambda_{0}\underbrace{\sum_{\ell}\tanh\!\big(C\,(\mathbf{z}_{\ell}\odot\|\mathbf{W}_{\text{dec}}^{\ell}\|)\big)}_{\text{$L_{0}$ sparsity}}$
$+\lambda_{\text{df}}\underbrace{\sum_{\ell}\mathrm{ReLU}\!\Big(\exp(\tau)-\mathbf{h}_{\ell}^{\text{pre}}\Big)\|\mathbf{W}_{\text{dec}}^{\ell}\|}_{\text{dead feature penalty}},$ (3)
where $\mathbf{W}_{\text{dec}}^{\ell}$ are concatenated decoder weights for layer $\ell$, $\mathbf{h}_{\ell}^{\text{pre}}$ are pre-activation values, $\tau$ is a threshold parameter, and $C$ is a scaling constant. $\lambda_{0}$ and $\lambda_{\text{df}}$ control the strength of the sparsity and dead-feature regularization terms.
We train CLTs on activation vectors sampled uniformly across languages until convergence. Training details, hyperparameters, and performance metrics appear in Appendix B.
3.3 Graph attribution
Following Ameisen et al. (2025), we compute the attribution score between every feature $n$ at layer $\ell$, position $k$, and feature $n^{\prime}$ at layer $\ell^{\prime}$, position $k^{\prime}$, as
$a_{\ell,k,n}^{\ell^{\prime},k^{\prime},n^{\prime}}=\sum_{\ell\leq s\leq\ell^{\prime}}f_{k,n}^{\ell\to s}\;J_{\ell,k}^{\ell^{\prime},k^{\prime}}\;g_{k^{\prime},n^{\prime}}^{\ell^{\prime}},$ (4)
where $f_{k,n}^{\ell\to s}$ denotes the vector for feature $n$ in the decoder matrix projecting from layer $\ell$ to $s$, $J_{\ell,k}^{\ell^{\prime},k^{\prime}}$ is the Jacobian between the MLP output at $(\ell,k)$ and the MLP input at $(\ell^{\prime},k^{\prime})$, computed during a forward pass where nonlinearities (normalization layers, attention computations, and MLPs) are frozen using stop-gradient operations, and $g_{k^{\prime},n^{\prime}}^{\ell^{\prime}}$ is the corresponding encoder feature at layer $\ell^{\prime}$ and position $k^{\prime}$. The graph is then pruned to retain only features that cumulatively account for $80\%$ of the effect on the final logit, and edges that cumulatively account for $95\%$ of the edge effect on the final logit. We use the circuit-tracer library Hanna et al. (2025).
3.4 Multilingual Score
For each CLT feature $f$, we compute the multilingual scores on both the top 100 most activated sequences for that feature and on a general sample of 600k sequences (each of length 16). If a feature is active for at least one token in a sequence, it is considered active for that sequence. Since each sequence corresponds to a specific language, this allows us to compute a normalized activation distribution:
$p_{l}(f)=\frac{A_{l}(f)}{\sum_{l^{\prime}}A_{l^{\prime}}(f)},$ (5)
where $A_{l}(f)$ is the number of sequences in language $l$ for which feature $f$ is active. We then define the multilingual score for feature $f$ as the entropy of this distribution:
$H(f)=-\sum_{l=1}^{L}p_{l}(f)\log p_{l}(f),\quad L=5,$ (6)
which measures multilinguality: low $H(f)$ indicates language-specific features, while high $H(f)$ indicates multilingual features.
4 Results
We organize our results around three core questions that progressively unpack how multilingual representations emerge and operate in large language models:
  1. 1.
    Do models form shared multilingual representations or a pivot language—across data mixtures?
  2. 2.
    How do models decode language identity and route information to the correct output language?
  3. 3.
    Why do models still fail in non-English settings despite shared representations?
4.1 Emergence of a Pivot Language
4.1.1 Generalization Under Data Imbalance
We begin by assessing whether models trained on imbalanced multilingual mixtures are still able to generalize across languages. Figure 2 reports the validation cross-entropy losses for models trained with varying proportions of English tokens, including extreme settings where English accounts for $90\%$ of the training data. Despite this imbalance, the models achieve competitive validation performance not only in English but also in minority languages such as Arabic, indicating that the learned representations capture transferable features rather than being restricted to high-resource languages. Together, these results suggest that multilingual circuits remain robust under imbalance and motivate a deeper investigation into how such robustness is organized within the model. More results are presented in Appendix E.
Figure 2: Validation cross-entropy loss curves for models trained on multilingual mixtures with varying proportions of English. Despite extreme imbalance (up to 90% English), the models maintain strong next-token prediction performance across languages.
4.1.2 Language Entropy Across Layers
To investigate how multilingual representations are distributed within the network, we analyze the language entropy of feature activations at each layer. For each feature, we compute the multilingual score $H(f)$ as defined in Section 3.4, and multiply it by its corresponding activation rate computed over a large subset of the training data. Figure 3 shows the resulting average weighted entropy for the top100 score across layers for models trained on different data mixtures.
Here we summarize three major findings on the layerwise organization of multilingual representations:
Finding 1.
Layerwise entropy follows a U-shaped trend. Entropy is relatively low in early layers, rises sharply in the middle layers, and decreases again toward the output layers. This indicates that early layers primarily encode language-specific features, middle layers integrate information into a shared multilingual space, and later layers partially re-specialize for language-specific decoding.
Finding 2.
The trend is robust across data mixtures and model scales. The emergence of a shared multilingual latent space in the middle layers is stable across different training data, and we also observe this well-defined behavior in the Llama 1B CLTs (see Figure 15), showing generalization to larger models.
Finding 3.
Model depth influences the behavior. For the 4-layer TinyStories model, the up-and-down pattern is absent and all layers have similar entropy scores, suggesting a minimum model size for this behavior to emerge.
The behavior is similar for the general entropy score. A discussion comparing the top100 score and the overall training distribution score is provided in Appendix D. These findings complement the validation loss results and provide quantitative evidence that the model organizes its internal representations to support cross-lingual generalization.
Figure 3: Rate-weighted multilingual score $H(f)$ across layers, showing that middle layers consistently form a multilingual space, while early and late layers are more language-specific. This pattern holds across model sizes and training data mixtures.
4.1.3 Multilingual Circuits Across Data Mixtures
To complement the multilingual score analysis, we extract attribution graphs to identify circuits active across data mixtures. We focus on two sentence types: (i) preposition sentences, where the model predicts a function word such as in J’ai bu une tasse or It was a piece, and (ii) content word sentences, where it predicts a meaningful token such as I prefer drinking tea to drinking in the 90% English mixture or Winter, spring, summer, and autumn are the four in the 20% mixture. We also analyze examples involving calendar terms (Monday, Tuesday, Wednesday, Thursday) and analogy prompts (the opposite of ’men’ is…). Figure 4 shows two representative circuits, one for an English preposition and one for a German content word, with remaining examples in Appendix F.
Figure 4: Circuits identified across different training data mixtures. Middle layers consistently form multilingual clusters, while early and late layers remain more language-specific. The scores at the top right of each cluster indicate language entropy (higher values correspond to more multilingual clusters). The patterns are stable across mixtures, sentence types, and languages.
Here we summarize three main findings on multilingual circuits across layers and data mixtures:
Finding 1.
Layerwise organization exhibits a three-phase structure: early and late layers are mostly language-specific, while middle layers form dense multilingual clusters. Circuits for determiners remain largely language-specific, confirming observations from Schut et al. (2025).
Finding 2.
Semantic alignment can emerge early or late depending on the data. In some cases in the 90% English model, the late-layer English cluster is missing, suggesting the model does not always rely on English for semantic grounding. Conversely, layer 0 sometimes already shows multilingual alignment. When tokenization is clean, the first attention and MLP blocks link semantically equivalent tokens across languages, indicating that semantic mapping begins from the very first layer. In the 20% mixture, Arabic features link words sharing historical meanings, such as connecting “qalb” (“heart”) to its older sense of “change” or “transformation,” showing that the model recovers deep etymological relations.
Finding 3.
Performance and representation quality are affected by tokenization. The model’s weaker performance on Arabic partially results from tokenization. Even with a balanced multilingual tokenizer, Arabic words are frequently split into small fragments, forcing the first layers (up to layer 3) to focus on reassembling words rather than learning higher-level meaning (Figures 21(b) and 22(b)).
Finding 4.
Feature alignment analysis shows no systematic bias toward English representations (Appendix L).
4.2 Mechanisms of Language Decoding
A central question in multilingual models with pivot languages is how models encode language information and integrate it in final layers to predict the correct next token. Across all graphs, language-specific features appear predominantly in early and late layers. In early layers, these features often correspond to single-token content rather than language alone. As in Lindsey et al. (2025b), we find early-layer features indicating output language for specific contexts (e.g., after quotation marks in “The opposite of ‘‘men’’ is ‘‘”), but these features are not consistently present in general cases.
Figure 5: Attribution graphs in French, English, and German for the 20% model pruned at 50% for the prompt “Autumn, Winter, Fall, and Spring are the four”. Interventions are displayed and language features highlighted in color with their corresponding percentage of activation on their language.
Figure 6: Frequency of token activations over their respective languages of the language features for the $20\%$ model.
Finding 1.
Across all attribution graphs, we find that late-layer language clusters drive decoding. These clusters correspond to high-frequency features that activate for large proportions of tokens within their respective languages. As shown in Appendix C, Figure 13, most high-frequency CLT features are language-specific. Figure 5 shows that these late-layer activations can be linearly read from early-layer features and embedding nodes, indicating that language identity propagates hierarchically through the model.
Finding 2.
Direct intervention experiments confirm the causal role of these clusters in controlling output language. Zeroing late-layer language features for one language and adding those from a target language (using activations from translated prompts) substantially increases the target-language logit rank. Perfect prediction typically requires additional feature modifications or sweeping intervention values.
Finding 3.
Finally, activation frequency patterns reveal both language ambiguity and dominance effects. As shown in Figure 6, a small number of features activate on 50–100% of tokens within each language, while inactive tokens occur mostly at sequence beginnings where language identity is ambiguous (especially between French, English, and German). Dominant languages exhibit fewer high-frequency features than others across both GPT-2 and LLaMA models, suggesting that English functions as a default output language requiring fewer explicit high-frequency features Lindsey et al. (2025b).
4.3 Failure Mechanisms in Non-English Languages
Our analysis in Section 4.1.3 showed that multilingual LLMs form shared representations in their middle layers, suggesting that equivalent meanings are processed through common internal mechanisms across languages. Yet, English-dominant models consistently perform better in English. If the model relies on shared circuits, why does this gap persist?
To answer this, we study two tasks probing distinct reasoning types:
  1. 1.
    Antonym Task: Prompts such as The opposite of ‘‘men’’ is ‘‘” test semantic relations.
  2. 2.
    Category Task: Prompts like “Football, cycling, baseball are all” probe conceptual grouping.
We analyze five languages (English, German, French, Arabic, Chinese) across four training mixtures (20%, 50%, 70%, 90% English dominance) to uncover the mechanisms behind non-English failures.
Figure 7: Circuit structures across language mixtures. Green: 90% English mixture, orange: 50% mixture. Key clusters missing in Arabic at 90% emerge by 50%, indicating mixture-dependent circuit formation. Additional examples appear in Appendix G.
4.3.1 Antonym Task
Circuits driving antonym prediction are consistent across languages and mixtures, but their completeness depends on mixture balance. In Arabic under the 90% English mixture, the Men&Women cluster, present in most other languages, is missing but appears from the 50% mixture onward (Figure 7). Intervention experiments (Figure 8) confirm its causal role: adding this missing cluster restores correct Arabic predictions (Appendix H).
We also find that the 20% mixture performs best for Latin-script languages (English, French, Arabic), as they share more transferable features than others (Appendix K).
Figure 8: Adding the missing Men&Women cluster to the Arabic circuit restores performance.
4.3.2 Category Completion Task
We analyze the Category Completion task using the 20% mixture to understand performance in the balanced setting. Arabic performs poorly due to tokenization fragmentation, which forces early layers to reconstruct words rather than activate answer-related clusters. Intervention experiments (Appendix I) show that scaling these clusters restores correct predictions, emphasizing the importance of activation strength.
Comparatively, English consistently triggers stronger responses in task-specific circuits (Men&Women for Antonym, Say Sports for Category Completion). This difference narrows under balanced mixtures but remains linked to sub-tokenization: languages with fragmented tokens show weaker edge strengths from embeddings to target clusters, partially explaining their reduced accuracy (Appendix J).
5 Implications
CLTs make it possible to trace how multilingual models process information step by step, from early encoding to shared representations and finally to language-specific decoding. By exposing feature interactions across layers, they reveal which parts of this pathway behave similarly across languages and where they start to diverge. This provides a concrete way to investigate language disparities: instead of relying only on task performance, we can directly observe which circuits are weaker, missing, or more fragmented for certain languages. More broadly, the ability to follow these internal mappings offers a structured foundation for future work on multilingual models, helping researchers identify which aspects of multilingual processing are stable and which are responsible for cross-language differences.
6 Conclusion
We present a mechanistic account of multilingual processing in LLMs using Cross-Layer Transcoders, showing that models form shared multilingual representations in their middle layers while relying on small sets of high-frequency features for language-specific decoding. These features linearly read out language identity from earlier layers, and targeted interventions on them can reliably shift the model’s output language. Performance differences across languages seem to arise less from missing multilingual circuits and more from factors such as tokenization and the strength of downstream activations. Overall, our results show how CLTs make it possible to pinpoint where multilingual disparities emerge and provide a practical basis for more focused investigations of cross-language processing in LLMs.
Limitations
Model Scale and Architecture. Our primary analysis focuses on relatively small models (68M and 177M parameters) trained from scratch. While we include LLaMA-1B for validation and observe consistent patterns, our findings may not fully generalize to production-scale models (10B+ parameters) or models trained with different architectural choices. Larger models may develop additional mechanisms or more nuanced multilingual representations that our current analysis does not capture.
Performance Disparity Analysis. Our investigation into why English outperforms other languages (Section 4.3) is based on a limited set of case studies across specific tasks (antonym and category completion). While these examples provide mechanistic insights into failure modes, they represent only a small sample of possible multilingual interactions. The findings, particularly regarding circuit activation strength and sub-tokenization effects, may not generalize to all task types or capture the full complexity of performance disparities across languages. A more comprehensive analysis would require systematic evaluation across diverse linguistic phenomena and larger example sets.
Tokenization Analysis. While we identify tokenization as a critical factor in performance disparities, our analysis uses a single BPE-based tokenizer trained on balanced data. We do not systematically compare alternative tokenization strategies (e.g., character-level, morpheme-aware, or language-specific tokenizers) that might better preserve semantic information across scripts. The optimal tokenization approach for multilingual models remains an open question.
Reproducibility Constraints. The computational cost of training both base models and CLTs (requiring thousands of GPU-hours) may limit independent replication. While we release our models and CLTs, the specific training dynamics and their sensitivity to random seeds, hardware variations, and implementation details have not been exhaustively characterized.
Ethical Considerations
Our analysis reveals systematic performance disparities across languages despite balanced training data. Arabic suffers from heavy sub-tokenization, forcing early layers to reconstruct fragmented words rather than process semantics, a structural disadvantage affecting billions of non-Latin script users. While we release our models and CLTs for reproducibility, training required thousands of GPU-hours, raising environmental concerns. Our intervention experiments demonstrate direct manipulation of model outputs across languages. Though intended for improving fairness, these techniques could potentially be misused. Our study covers only five languages and small models (68M, 177M and 1B parameters), limiting generalizability to the world’s  7,000 languages and production-scale systems. We hope these mechanistic insights contribute to more equitable multilingual AI development.
Acknowledgment
We thank Michael Hanna for insightful discussions on the added value of Cross-Layer Transcoders compared to Single-Layer Transcoders, as well as on the training challenges and implementation details of CLTs. We are grateful to Anson Lei for his assistance with the trained GPT-2 model generation issue and for sharing tips on prompt formatting. We also thank Francesco Ortu for his valuable feedback, which helped improve the clarity and flow of the paper.
This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by Schmidt Sciences SAFE-AI Grant; by the Frontier Model Forum and AI Safety Fund; by Open Philanthropy; by the Cooperative AI Foundation; and by the Survival and Flourishing Fund. Resources used in preparing this research project were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. F. D. acknowledges support through a fellowship from the Hector Fellow Academy.
Appendix A Multilingual Model Training
A.1 Tokenization Strategy
All models employ a unified tokenization approach designed to eliminate tokenization-induced biases. We train a debiased BPE tokenizer on uniformly distributed five-language data, with each language contributing exactly the same number of tokens. The resulting vocabulary contains 119,547 tokens, ensuring balanced representation across languages and writing systems.
Tokenizer Training Protocol
  • Training Data: Equal token counts per language (20% each)
  • Vocabulary Size: 119,547 tokens
  • Algorithm: Byte-Pair Encoding (BPE)
  • Special Tokens: BOS (beginning of sentence), EOS (end of sentence), PAD (padding), UNK (unknown)
Architecture Layers Hidden Dim Attention Heads Parameters Context Length
GPT-2 12 768 12 177.6M 1024
TinyStories 4 768 16 68.5M 512
Table 1: Model architecture specifications.
A.2 Model Architectures and Training
A.2.1 Model Architecture Specifications
We train transformer models following two architectures to enable comparisons across model families and scales:
Component GPT-2 TinyStories
Vocabulary Size 119,547 119,547
Embedding Dimension 768 768
FFN Inner Dimension 3072 3072
Attention Head Dim. 64 48
Dropout Rate 0.1 0.1
Layer Norm Pre-LN Pre-LN
Activation Function GELU GELU
Position Encoding Learned Learned
Context Length 1024 512
Table 2: Detailed architecture hyperparameters.
A.2.2 Training Scale and Compute Optimization
Following Chinchilla scaling laws Hoffmann et al. (2022), we compute optimal training data sizes:
$N_{\text{tokens}}=20\times N_{\text{parameters}}$ (7)
Model Tokens Steps FLOPs
GPT-2 3.55B 1.73M $3.8\times 10^{12}$
TinyStories 1.37B 670K $5.6\times 10^{11}$
Table 3: Compute-optimal training specs.
A.2.3 Training Implementation
We use the nanoGPT codebase (Karpathy, 2022) for reproducible transformer training.
Optimization Configuration
Parameter Value
Optimizer AdamW
Learning Rate 6e-4
$\beta_{1}$ 0.9
$\beta_{2}$ 0.95
Weight Decay 0.1
Gradient Clipping 1.0
Warmup Steps 2000
LR Schedule Cosine linear warmup
Batch Size 64
Gradient Accumulation 8 steps
Effective Batch Size 512
Table 4: Training hyperparameters.
Training Infrastructure
  • Hardware: 4×H100 80GB GPUs per model
  • Distributed Training: Data parallel with gradient synchronization
  • Mixed Precision: FP16 training with automatic loss scaling
  • Checkpointing: Every 5000 steps with automatic resumption
  • Monitoring: Real-time loss and gradient norm tracking
A.2.4 Training Dynamics
Figure 2 shows training and validation loss across English data proportions. Higher English proportions reduce English validation loss but can slightly degrade performance on other languages.
A.2.5 Training Procedure and Model Matrix
Each architecture is trained on all data mixtures, producing a systematic model matrix:
Architecture 20% 50% 70% 90%
GPT-2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TinyStories $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Total Models: 8
Table 5: Complete model training matrix.
For comparative purposes, we also include a pre-trained LLaMA-1B model alongside our custom-trained models.
Appendix B CLTs training
B.1 Training Configuration
We report in Table 6, the hyper-parameters of the CLT training. We mostly follow Anthropic guidelines Ameisen et al. (2025). This setup is similar across Tinystories, GPT-2, and Llama CLTs.
Table 6: Training Configuration
Parameter Value
Model Architecture
Input dimension ($d_{\text{in}}$) 768
Latent dimension ($d_{\text{latent}}$) 24,576
Expansion factor 32
Context size 16
Training Hyperparameters
Learning rate $2\times 10^{-4}$
Adam $\beta_{1}$ / $\beta_{2}$ 0.9 / 0.999
Batch size (tokens) 1,024
LR warm-up steps 1,000
LR decay steps 3,749
Loss Coefficients
L0 coefficient 2.0
Optimal L0 10
Dead penalty coefficient $1\times 10^{-5}$
JumpReLU Configuration
Bandwidth 1.0
Initial threshold 0.03
Other
Seed 42
Precision float32 (mixed)
However, for the Llama CLT, we use less features, as we are limited by GPU size with an expansion factor of around 4. We use torch.float32 for the training as well.
B.2 Performance
We report the amount of dead features and the performance of the different CLTs in Figures  9  10  11.
Figure 9: Dead feature count and explained variance across layers for the GPT-2 CLTs.
Figure 10: Explained variance over layers for the Llama CLT.
Figure 11: Dead feature count and explained variance across layers for the Tinystories CLTs.
Appendix C Language Features
We add here two plots related to the language features.
Figure 12: Activation rate over the top100 sequences of the language feature for the 20% model. We see that some language features are more active towards the end of the sequence. Highlighted is the language feature with the highest frequency.
Figure 13: CLT features with the token activation frequency above 5% for the 20% model vs the probability in their top activating language. This shows that most high-frequency features are language features.
Appendix D Multilingual Score
We compare the top100 multilingual score with the multilingual score computed over a large training subset (Figure 14). Both entropy measures show the same overall trend across layers, with the top100 entropy consistently shifted toward lower values. The main difference appears in the earliest layers. This is likely because many features in these layers correspond to single-token activations, especially in layer 0. Given that the total number of features is limited to approximately 28K and multiple languages are involved, some features that mostly activate for a single token also tend to activate for other unrelated tokens.
Figure 14 also shows the frequency of the most activated token for each feature across the top100 sentences where that feature is strongest. We observe that in layer 0, and to a lesser extent in layers 1 and 2, most activations correspond to single tokens. For this reason, we decide to report the top100 score on the graphs.
Figure 14: Comparaison between the two multilingual entropy measures. One computed on the top100 activated sequences for each feature and one computed over a large training subset. On the right the frequency of activation of the most frequent token over the top100 sequences.
We also report the general multilingual scores for Tinystories and Llama 3.2-1B in Figure 15.
Figure 15: General multilingual score for the Tinystories model and Llama 3.2-1B. The final layer jump in the Llama CLT is potentially due to training dynamics. We find that the last layer has some features activated with small values that impact slightly the reconstruction performance but have a great impact on the multilingual score.
During graph analysis, we display the multilingual score and the language distributions on a graph visual interface as shown in Figure 16.
Figure 16: Graph visual interface with multilingual distributions and entropies on the top right.
Appendix E Additional Validation Loss Curves Under Data Imbalance
In Section 4.1.1, we reported validation loss curves for four representative languages (English, Arabic, German, French). Here, we provide additional results for (i) Chinese and (ii) models trained on the TinyStories dataset.
E.1 Chinese
Figure 17 shows the validation loss curves for Chinese. The results mirror the patterns observed for other languages: despite extreme imbalance in the training data, the model maintains competitive prediction performance in Chinese, indicating that the robustness of multilingual circuits extends to typologically distinct languages.
Figure 17: Validation cross-entropy loss curves for Chinese under varying proportions of English. The same robustness pattern holds as for other languages.
E.2 TinyStories Models
To assess whether the observed generalization dynamics also hold at small scale, we trained models on the TinyStories dataset. Figure 18 reports validation loss curves for five languages (English, Arabic, German, French, Chinese). Despite the reduced data scale and synthetic nature of TinyStories, the same trend emerges: performance remains robust across all languages, confirming that the organization of multilingual circuits is a stable property that persists across both large-scale and small-scale training regimes.
Figure 18: Validation cross-entropy loss curves for the TinyStories-trained models across five languages (English, Arabic, German, French, Chinese). Despite small-scale training data, models retain strong multilingual generalization under imbalance.
Appendix F Additional Results: Multilingual Circuits
In this appendix, we provide extended circuit visualizations complementing the analysis of Section 4.1.3. For each example, we extract attribution graphs across layers and present the identified circuits. Depending on the setting, each figure contains either (i) four subfigures, corresponding to the four training data mixtures (90%, 70%, 50% or 20% English), or (ii) five subfigures, corresponding to the five languages (English, Arabic, German, French, Chinese).
As observed in our main results, early and late layers predominantly encode language-specific circuits, while middle layers form clusters with high language entropy, reflecting multilingual representations. Exceptions arise for determinants and prepositions, whose circuits remain largely language-specific.
We now detail the examples analyzed.
F.1 Preposition Sentences (English: ‘‘It was a piece’’)
We begin by analyzing English preposition prediction across training mixtures. Prepositions, being primarily functional rather than semantic, offer insight into the model’s handling of syntactic structures. In this case, we find that the model consistently relies on language-specific clusters.
(a) 90% mixture
(b) 70% mixture
(c) 50% English mixture
(d) 20% multilingual mixture
Figure 19: Circuits for English preposition prediction (“It was a piece”) across training mixtures.
F.2 Preposition Sentences (French: ‘‘J’ai bu une tasse’’)
We next examine French preposition prediction across the training mixtures. This setting tests whether the patterns observed for English prepositions generalize to another language.
(a) 90% mixture
(b) 70% mixture
(c) 20% mixture
(d) 70% mixture
Figure 20: Circuits for French preposition prediction (“J’ai bu une tasse”) across training mixtures.
F.3 Content Word Sentences (English: ‘‘I prefer drinking tea to drinking’’)
Next, we analyze content word prediction across languages in the 90% English mixture. Content words engage semantic and contextual representations, making them a strong testbed for cross-lingual generalization. We find robust multilingual clusters in middle layers across all five languages, even when trained under extreme imbalance.
(a) English
(b) Arabic
(c) German
(d) French
(e) Chinese
Figure 21: Circuits for content word prediction (“I prefer drinking tea to drinking”) across languages in the 90% English mixture.
F.4 Content Word Sentences: ‘‘Winter, spring, summer, and autumn are the four’’
We analyze the prediction of content words in the sentence “Winter, spring, summer, and autumn are the four” across languages in the 20% English mixture. This example probes semantic and sequential reasoning across multiple languages, highlighting whether the model reuses internal circuits for similar content words. We find that middle layers consistently form multilingual clusters across all five languages, while early and late layers remain largely language-specific.
(a) English
(b) Arabic
(c) German
(d) French
(e) Chinese
Figure 22: Circuits for content word prediction (“Winter, spring, summer, and autumn are the four”) across five languages in the 20% English mixture.
F.5 Calendar Term Prediction (‘‘Monday, Tuesday, Wednesday, Thursday’’)
We study calendar term prediction across mixtures. Calendar terms primarily involve lexical memorization, providing a probe for multilingual circuit reuse. Across all mixtures, middle layers consistently form multilingual clusters, while early and late layers remain more language-specific.
(a) 90% mixture
(b) 70% mixture
(c) 50% mixture
(d) 20% mixture
Figure 23: Circuits for calendar term prediction (“Monday, Tuesday, Wednesday, Thursday”) across mixtures.
F.6 Analogy Sentences (‘‘the opposite of ’men’ is ...’’)
Finally, we analyze analogy-based prediction across mixtures. Unlike lexical recall, analogies engage relational reasoning. We find that the same multilingual clusters are reused across mixtures, especially in middle layers, underscoring their role in cross-lingual generalization.
(a) English
(b) French
(c) German
(d) Arabic
(e) Chinese
Figure 24: Circuits for analogy-based prediction (“the opposite of ’men’ is …”) across mixtures.
Appendix G Why Do Models Fail in Non-English Languages? Complementary Circuits
We provide comprehensive circuit visualizations demonstrating how circuit structures differ across languages and training mixtures, revealing the mechanisms behind performance disparities discussed in Section 4.3.
G.1 Circuit Analysis Across Mixtures
G.1.1 90% English Mixture
Figure 25 shows circuit graphs for the antonym prompt “The opposite of ‘‘men’’ is ‘‘” across all five languages.
(a) English
(b) French
(c) German
(d) Arabic
(e) Chinese
Figure 25: Circuits identified across five languages at 90% mixture.
The circuits reveal why English achieves superior performance: languages that fail to predict the correct antonym consistently lack the Men&Woman cluster. Arabic shows particularly sparse connectivity, missing most multilingual clusters found in other languages due to limited circuit sharing with English.
G.1.2 70% English Mixture
(a) English
(b) French
(c) German
(d) Arabic
Figure 26: Circuits identified across five languages at 70% mixture.
Figure 26 shows identical patterns: the Men&Women cluster remains absent where models fail. Only English sufficiently activates the reasoning circuit at this mixture level.
G.1.3 50% English Mixture
(a) English
(b) French
(c) German
(d) Chinese
Figure 27: Circuits identified across five languages at 50% mixture.
Figure 27 validates previous patterns while showing increased presence of Men&Women clusters in non-English languages, reflecting the shift toward more uniform language distribution.
G.1.4 20% English Mixture
(a) English
(b) French
(c) Arabic
(d) Chinese
Figure 28: Circuits identified across five languages at 20% mixture.
Figure 28 shows more Men&Women clusters and improved performance, especially for Latin-script languages. Even at this balanced mixture, non-Latin script languages still demonstrate weaker circuit formation.
G.2 Competing Circuit Mechanisms
The circuit patterns suggest two competing mechanisms: a reasoning circuit that correctly predicts Women and a copying circuit that simply reproduces Men from the input tokens.
To validate these hypotheses, we conduct four investigations:
  1. 1.
    Intervention experiments: scale up reasoning circuit activation (Men&Women cluster) while scaling down copying circuit activation (Men cluster)
  2. 2.
    Comparative analysis across the three remaining mixtures to identify general trends
  3. 3.
    Activation value analysis of reasoning and copying circuits to validate insufficient activation hypothesis
  4. 4.
    Investigation of why English, German, and French demonstrate superior performance across all mixtures
Appendix H Intervention Validation
Figure 29 demonstrates successful interventions that validate the competing circuits hypothesis.
(a) 20% mixture model intervention
(b) 50% mixture model intervention
(c) 70% mixture model intervention
(d) 90% mixture model intervention
Figure 29: Intervention examples where appropriate scaling switches the model’s answer to the correct response.
In all the examples where the model fails to get the right answer, scaling up the reasoning circuit while scaling down the copying circuit consistently produces correct predictions. We sweep positive coefficients for the reasoning circuit (1 to 30) and negative coefficients for the copying circuit (-30 to -1) to determine optimal steering values.
Figures  28,  27,  26, and  25 further validate our hypothesis. In particular, most of the examples where the model fails to predict the correct answer show either an absence of the reasoning circuit or a low level of its activation, and this pattern becomes less pronounced as we move towards the 20% mixture. Taken together with the intervention results, this supports our hypothesis about circuit competition.
Appendix I Category Completion Task Circuits and Interventions
Here, we present the circuits of the Category Completion Task presented in Section 4.3.2.
(a) German
(b) French
(c) Arabic
(d) Chinese
Figure 30: Circuits for category completion task.
We also present the intervention results for Arabic in Figure 31, the only language where the model failed.
Figure 31: Scaling up the ‘‘Say Sports’’ cluster makes the model predict the right answer.
Appendix J Why English Performs Better: A Case Study
We present a case study examining the 20% mixture to understand performance disparities in balanced training.
J.1 Cluster Activation Analysis
We measure activation strength of task-relevant clusters across languages. For the Antonym task, we track the Man&Woman cluster; for Category Completion, the Say Sports cluster.
Figure 32 shows activation patterns. English consistently produces higher activations than other languages. This disparity decreases approaching the 20% mixture, suggesting balanced training mitigates but does not eliminate the advantage.
Figure 32: Cluster activation strength across languages and mixture ratios. English shows consistently higher activations for task-relevant clusters (Man&Woman for Antonym, Say Sports for Category Completion), with disparities diminishing at the 20% mixture.
J.2 Sub-tokenization Effects
We quantify sub-tokenization impact by measuring edge strength from token embeddings to target clusters. Edge strength is the dot product between the residual stream after token embedding and the cluster’s input direction.
Figure 33 presents results across languages with varying sub-tokenization rates. Highly sub-tokenized languages (e.g., Arabic) show significantly weaker edge strengths. Fragmented tokens fail to properly activate downstream circuits—semantic information distributes across multiple sub-tokens rather than concentrating in a single embedding.
Figure 33: Edge strength from token embeddings to target clusters versus sub-tokenization rate. Higher sub-tokenization correlates with weaker edge strength (r=-0.82), explaining reduced circuit activation in fragmented languages like Arabic.
Appendix K Latin Script Performance Advantage
Figure 34: Feature overlap between each language and English in the 20% multilingual model.
Figure 34 demonstrates that English shares more features with French and German than with Arabic and Chinese. This feature overlap explains the superior performance of Latin-script languages through shared circuit structures, even at the most balanced mixture.
Appendix L Feature Alignment with English Analysis
L.1 Methodology
We investigate whether multilingual circuits preferentially align with English representations by comparing feature alignment with native versus English tokens across layers.
For each example, we extract features contributing to the correct answer using circuit tracing. For each feature at layer $\ell$, we compute its decoder vector $\mathbf{d}_{\ell}$ and measure alignment with the unembedding space:
$\text{sim}(t)=\frac{\mathbf{d}_{\ell}}{\|\mathbf{d}_{\ell}\|}\cdot\mathbf{W}_{U}[:,t]$ (8)
where $\mathbf{W}_{U}$ is the unembedding matrix and $t$ is a token. We compute similarity for both the native answer token and its English equivalent, then rank all vocabulary tokens by similarity. Lower ranks indicate stronger alignment.
We normalize ranks to $[0,1]$ via $\text{norm}(r)=1-\frac{r-r_{\min}}{r_{\max}-r_{\min}}$, where higher values indicate better alignment. We aggregate results across layers and languages.
L.2 Results
Figure 35 shows average alignment across all languages. Both native and English tokens show similar ranking patterns across layers, with no systematic bias toward English.
(a) 20% mixture
(b) 50% mixture
(c) 70% English mixture
(d) 20% multilingual mixture
Figure 35: Feature alignment with native versus English tokens across mixture ratios. Normalized ranks show no systematic bias toward English representations across all training mixtures, with native and English tokens exhibiting similar alignment patterns throughout the network depth.